1. The propeller of a World War II fighter plane is 2.38 m in diameter.

a.  What is its angular velocity in radians per second if it spins at 1100 rev/min?

b. What is the linear speed (in m/s) of its tip at this angular velocity if the plane is stationary on the tarmac?

c.  What is the centripetal acceleration of the propeller tip under these conditions? Calculate it in meters per second squared and convert to multiples of g.

centripetal acceleration in m/s2 =?

  centripetal acceleration in g =?
Answer #1

1. Given,

The diameter, d = 2.30 m

So the radius, r = 2.30 / 2 = 1.15 m

a)

Given, = 1100 rev/min

= 1100 * (2/60) = 115.17 rad s

b)

The linear velocity, v = * r

v = 115.17 * 1.15

= 132.45 m/s

c)

The centripetal acceleration, a = 2 r

= (115.17)2 * 1.15

= 15253.75 m/s2

The centripetal acceleration in terms of g, a = 1557 g